# COMAE: COMprehensive Attribute Exploration for Zero-shot Hashing

### v1

@@ 4.1 Experimental Settings
Figure 3: mAP@all and AUC scores and their decline percentage with the change of unseen classes ratio in the training process.
4.1.1 Datasets
In our investigations, we meticulously analyzed three distinguished zero-shot hashing datasets referenced in [32], i.e., AWA2 [43], CUB [39], SUN [26]. AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39]. The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes. SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
4.1.2 Baselines
@@
  • Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH [1], SH [42], ITQ [10], SDH [29], IMH [30], and PCA [25]. Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
  • Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images. Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet [3], GreedyHash [35], JMLH [31], ADSH [52], CSQ [48], DPN [7], OrthoCos [14], and unsupervised hashing works, i.e, BiHalf [20], TBH [28], CIBHash [28]. By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
  • Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training. These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes. Representative works include TSK [46], ASZH, SitNet, OPZH, and AH. ASZH [32] uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data. SitNet [11], and OPZH [49] adopt semantic vectors to find semantic embedding space for knowledge transferability. AH [45] constructs relationships and image features for transferability.
4.1.3 Evaluation Metric
@@
4.1.4 Implementation Details
All comparative baselines are conducted with their default hyper-parameter configurations. For a fair comparison, their image feature extractors are substituted with the ResNet101 [17]. The training epoch is set to 10 and the batch size is set to 64, and more than 10 will be overfitting, resulting in a decrease in accuracy. We take Adam as the optimizer weight decay 0.0005 and the learning rate is set to 0.0001. The $\epsilon$ in Eqn. 9 is set to 0.9, the $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ in Eqn. 4 are set to 10, 1, and 10 respectively. To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
@@ 4.1.4 Implementation Details
Figure 4: The comparison of PR Curve, P@N Curve and R@N Curve in 64 bit length codes.
4.2 RQ1: Overall Performance
Table. 1 displays the results of the COMAE and all competitive baselines on datasets AWA2, CUB, and SUN with hash code lengths of 24, 48, 64, and 128. Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in paper [32]. The experimental settings of Table 1 are the same as [32]. In Fig. 4, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets. Typically, the efficacy of a method is gauged by the PR curve approaching the upper right quadrant, with higher positions on the P@N and R@N curves indicative of superior performance. Overall, we find the following conclusions,
@@ 4.2 RQ1: Overall Performance
  • COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths. At the same time, the performance of COMAE remains relatively stable when the number of bits increases. Compared with AH [45], an attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth.
@@
4.3 RQ1: Hashing Code Visualization
In this experiment, we adopted a more intuitive approach to visualizing hashing codes to evaluate the effectiveness of our proposed COMAE. Throughout the comparative experiments, we initially transformed the 64-bit hashing codes into lower-dimensional spaces, subsequently projecting these codes into a 2-D plane using the t-SNE [24] tool. This methodology facilitated a clearer observation of COMAE’s performance through visual means. We present the visualization of hash code for all testing samples with unseen classes on the AWA2 dataset. Results are shown in Fig.  7. According to the visualization results, we find that a) HashNet, DPN and OrthoCos perform badly, which are unable to distinguish the unseen classes. b) Although CSQ and BiHalf gain favorable outcomes in visualization and mAP@5000 results in Table. 1, they still encounter challenges such as substantial inter-class overlap and indistinct boundaries (the classes represented by the green and other colors). c) Compared with other baselines, the proposed COMAE achieves the best performance, as the generated hash codes are more distinguished. It indicates that COMAE can effectively classify images.
@@ 4.3 RQ1: Hashing Code Visualization
Figure 5: Histogram distances of the intra-classes and inter-classes. The arrow annotation is the quantitative separability with the hamming distance, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$.
4.4 RQ2: Effects of Unseen Class Ratio
To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories underwent a progressive reduction during the training process. The results are visually depicted in Fig.  3. Specifically, (d)-(f) and (j)-(l) show the performance decrease ratio between 20%->40% and 20%->60% respectively. We conclude the following conclusions based on the outcome results,
@@ 4.4 RQ2: Effects of Unseen Class Ratio
Figure 6: The effects of hyper-parameters on AWA2 dataset.
@@
Figure 7: t-SNE visualizations of hashing code on test unseen classes. The 10 colors denote 10 different unseen classes.
4.5 RQ3: Distance Visualization of Intra and Inter Classes
@@ 4.6 RQ3: Component Analysis
Table 2: The mAP@all and AUC results with switching model components in different datasets.
@@ 5 Conclusion
Acknowledgments
This paper was supported by the Natural Science Foundation of China under Grant No. 61836013 and the Informatization Plan of Chinese Academy of Sciences CAS-WX2023ZX01-1103.

### v2

@@ 3.1. Experimental Settings
Figure 3. mAP@all and AUC scores and their decline percentage with the change of unseen classes ratio in the training process.
3.1.1. Datasets
In our investigations, we meticulously analyzed three distinguished zero-shot hashing datasets referenced in (Shi et al., 2022), i.e., AWA2 (Xian et al., 2018), CUB (Wah et al., 2011), SUN (Patterson and Hays, 2012). AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39]. The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes. SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
3.1.2. Baselines
@@
  • Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008), SH (Weiss et al., 2008), ITQ (Gong et al., 2012), SDH (Shen et al., 2015a), IMH (Shen et al., 2015b), and PCA (Maćkiewicz and Ratajczak, 1993). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
  • Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images. Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet (Cao et al., 2017), GreedyHash (Su et al., 2018), JMLH (Shen et al., 2019), ADSH (Zhou et al., 2019), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), OrthoCos (Hoe et al., 2021), and unsupervised hashing works, i.e, BiHalf (Li and van Gemert, 2021), TBH (Qiu et al., 2021), CIBHash (Qiu et al., 2021). By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
  • Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training. These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes. Representative works include TSK (Yang et al., 2016), ASZH, SitNet, OPZH, and AH. ASZH (Shi et al., 2022) uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data. SitNet (Guo et al., 2017), and OPZH (Zhang et al., 2019) adopt semantic vectors to find semantic embedding space for knowledge transferability. AH (Xu et al., 2017) constructs relationships and image features for transferability.
3.1.3. Evaluation Metric
@@
3.1.4. Implementation Details
All comparative baselines are conducted with their default hyper-parameter configurations. For a fair comparison, their image feature extractors are substituted with the ResNet101 (Khan et al., 2018). The training epoch is set to 10 and the batch size is set to 64, and more than 10 will be overfitting, resulting in a decrease in accuracy. We take Adam as the optimizer weight decay 0.0005 and the learning rate is set to 0.0001. The $\epsilon$ in Eqn. 9 is set to 0.9, the $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ in Eqn. 4 are set to 10, 1, and 10 respectively. To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
@@ 3.1.4. Implementation Details
Figure 4. The comparison of PR Curve, P@N Curve and R@N Curve in 64 bit length codes.
3.2. RQ1: Overall Performance
Table. 1 displays the results of the COMAE and all competitive baselines on datasets AWA2, CUB, and SUN with hash code lengths of 24, 48, 64, and 128. Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in paper (Shi et al., 2022). The experimental settings of Table 1 are the same as (Shi et al., 2022). In Fig. 4, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets. Typically, the efficacy of a method is gauged by the PR curve approaching the upper right quadrant, with higher positions on the P@N and R@N curves indicative of superior performance. Overall, we find the following conclusions,
@@ 3.2. RQ1: Overall Performance
  • COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths. At the same time, the performance of COMAE remains relatively stable when the number of bits increases. Compared with AH (Xu et al., 2017), an attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth.
@@
3.3. RQ1: Hashing Code Visualization
In this experiment, we adopted a more intuitive approach to visualizing hashing codes to evaluate the effectiveness of our proposed COMAE. Throughout the comparative experiments, we initially transformed the 64-bit hashing codes into lower-dimensional spaces, subsequently projecting these codes into a 2-D plane using the t-SNE (Van der Maaten and Hinton, 2008) tool. This methodology facilitated a clearer observation of COMAE’s performance through visual means. We present the visualization of hash code for all testing samples with unseen classes on the AWA2 dataset. Results are shown in Fig.  7. According to the visualization results, we find that a) HashNet, DPN and OrthoCos perform badly, which are unable to distinguish the unseen classes. b) Although CSQ and BiHalf gain favorable outcomes in visualization and mAP@5000 results in Table. 1, they still encounter challenges such as substantial inter-class overlap and indistinct boundaries (the classes represented by the green and other colors). c) Compared with other baselines, the proposed COMAE achieves the best performance, as the generated hash codes are more distinguished. It indicates that COMAE can effectively classify images.
@@ 3.3. RQ1: Hashing Code Visualization
Figure 5. Histogram distances of the intra-classes and inter-classes. The arrow annotation is the quantitative separability with the hamming distance, $\mathbb{E}[D_{inter}]-\mathbb{E}[D_{intra}]$.
3.4. RQ2: Effects of Unseen Class Ratio
To verify the robustness of our model in the context of the zero-shot scenario, we conduct experiments where the number of seen class categories underwent a progressive reduction during the training process. The results are visually depicted in Fig.  3. Specifically, (d)-(f) and (j)-(l) show the performance decrease ratio between 20%-¿40% and 20%-¿60% respectively. We conclude the following conclusions based on the outcome results,
@@ 3.4. RQ2: Effects of Unseen Class Ratio
Figure 6. The effects of hyper-parameters on AWA2 dataset.
@@
Figure 7. t-SNE visualizations of hashing code on test unseen classes. The 10 colors denote 10 different unseen classes.
3.5. RQ3: Distance Visualization of Intra and Inter Classes
@@ 3.6. RQ3: Component Analysis
Table 2. The mAP@all and AUC results with switching model components in different datasets.